OpenStudy (anonymous):

2x+14y=8 solve for x

OpenStudy (anonymous):

what am i suppose to do

OpenStudy (midhun.madhu1987):

Subtract 14y from both sides.. first... what u get?

OpenStudy (anonymous):

give me steps then ill solve

OpenStudy (midhun.madhu1987):

and then divide by2 and u get the value of x.

OpenStudy (igreen):

2x + 14y = 8 Subtract 14y to both sides: 2x = 8 - 14y Divide to both sides of the equation: x = 4 - 7y

OpenStudy (anonymous):

oh wow so simple i thought i had to do something with the y intercept form or something